# Dual Quaternion SE(3) Synchronization with Recovery Guarantees

### v1

@@ 3 Iterative Estimation Performance
It is worth noting the improvement in estimation accuracy provided by DQGPM over the rounded spectral estimator. Theorem 3.2 establishes that, asymptotically as $k\to\infty$ (assuming a small noise constant regime where $\|\bm{\Delta}\|_{\mathrm{op},st},\|\bm{\Delta}\|_{\mathrm{op},\mathcal{I}}\leq an$ for sufficiently small $a>0$), the DQGPM estimation error satisfies:
$\operatorname{d}_{{}_{st}}(\bm{x}^{\infty},\hat{\bm{x}})\lesssim\frac{12}{n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,st},\ \operatorname{d}_{\mathcal{I}}(\bm{x}^{\infty},\hat{\bm{x}})\lesssim\frac{12}{n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,\mathcal{I}}.$
In contrast, Theorem 2.8 guarantees an error bound for the rounded spectral estimator of order $8\|\bm{\Delta}\|_{\mathrm{op}}/\sqrt{n}$. Comparing these rates, DQGPM yields a tighter bound provided that $\|\bm{\Delta}\hat{\bm{x}}\|_{2,st}\leq c\sqrt{n}\|\bm{\Delta}\|_{\mathrm{op},st}$ and $\|\bm{\Delta}\hat{\bm{x}}\|_{2,\mathcal{I}}\leq c\sqrt{n}\|\bm{\Delta}\|_{\mathrm{op},\mathcal{I}}$ for a constant $c<2/3$. Note that the trivial bound $c=1$ holds automatically by definition. In many stochastic noise models, however, $\hat{\bm{x}}$ is not aligned with the top singular directions of $\bm{\Delta}$, so $\|\bm{\Delta}\hat{\bm{x}}\|_{2}$ is typically much smaller than the worst-case bound (e.g., $c\approx 1/2$ for standard i.i.d. Gaussian noise), which highlights the advantage of the DQGPM refinement.
@@
Suppose the noise assumptions in Theorem 3.2 hold. Let $\bm{C}\in\mathbb{DH}^{n\times n}$ be Hermitian and admits a unique dominant eigenvalue with multiplicity one. Denote its right eigenvalues by $\lambda_{1}>\lambda_{2}\geq\cdots\geq\lambda_{n}$, where each $\lambda_{i}=\lambda_{i,\mathrm{st}}+\lambda_{i,\mathcal{I}}\epsilon$ is a dual number. If we run Algorithm 1 for $K_{\mathrm{init}}$ iterations satisfying
$K_{\mathrm{init}}\geq\max\Biggl\{\log_{r}\left({\frac{35|\alpha_{2,st}|+34M_{st}}{|\alpha_{1,st}|}}\right),\log_{r}\left({\frac{25N_{2}+24M_{\mathcal{I}}}{N_{1}}}\right)\Biggr\},$
then, initializing Algorithm 2 with $\bm{x}^{0}:=\sqrt{n}\,\bm{w}^{K_{\mathrm{init}}}$, the resulting sequence $\{\bm{x}^{k}\}_{k\geq 0}$ satisfies:
[…]
and
$\operatorname{d}_{\mathcal{I}}(\bm{x}^{k},\hat{\bm{x}})\leq\left(\frac{1}{10}\right)^{k}\frac{\sqrt{n}}{18}+k\left(\frac{1}{10}\right)^{k-1}\frac{\sqrt{n}}{200}+\frac{700}{53n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,\mathcal{I}}+\frac{875}{477n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,st}.$
Here $r:=\bigl|\lambda_{1,_{st}}/\lambda_{2,_{st}}\bigr|>1$, $M_{{}_{st}},M_{\mathcal{I}},N_{1},N_{2}$ are some real constants depending on $\alpha_{1},\lambda_{1},\lambda_{2}$, and $\alpha_{1}$ is the projection of initializer $\bm{w}^{0}$ onto normalized dominant eigenvector.
@@ 4 Numerical Experiments
Table 2: Rotation and translation errors of DQGPM and EIG across noise levels and observation rates.
$p=0.05$ Error_r DQGPM 0.034 $\pm$ 0.051 0.035 $\pm$ 0.051 0.036 $\pm$ 0.050 0.037 $\pm$ 0.049
EIG 0.056 $\pm$ 0.632 0.050 $\pm$ 0.183 0.052 $\pm$ 0.130 0.058 $\pm$ 0.170
Error_t DQGPM 0.069 $\pm$ 0.043 0.132 $\pm$ 0.064 0.196 $\pm$ 0.088 0.260 $\pm$ 0.112
EIG 0.580 $\pm$ 1.924 0.662 $\pm$ 1.982 0.739 $\pm$ 5.556 0.807 $\pm$ 2.397
$p=0.08$ Error_r DQGPM 0.001 $\pm$ 0.008 0.002 $\pm$ 0.008 0.003 $\pm$ 0.008 0.004 $\pm$ 0.008
EIG 0.001 $\pm$ 0.625 0.003 $\pm$ 0.013 0.004 $\pm$ 0.006 0.005 $\pm$ 0.878
Error_t DQGPM 0.032 $\pm$ 0.007 0.064 $\pm$ 0.009 0.095 $\pm$ 0.012 0.127 $\pm$ 0.014
EIG 0.042 $\pm$ 0.676 0.085 $\pm$ 0.814 0.128 $\pm$ 0.922 0.170 $\pm$ 1.062
$p=0.30$ Error_r DQGPM 0.0005 $\pm$ 0.000 0.001 $\pm$ 0.000 0.001 $\pm$ 0.000 0.002 $\pm$ 0.000
EIG 0.001 $\pm$ 0.000 0.001 $\pm$ 0.000 0.002 $\pm$ 0.625 0.002 $\pm$ 0.000
Error_t DQGPM 0.013 $\pm$ 0.001 0.027 $\pm$ 0.001 0.040 $\pm$ 0.002 0.053 $\pm$ 0.003
EIG 0.015 $\pm$ 0.033 0.031 $\pm$ 0.130 0.047 $\pm$ 0.113 0.062 $\pm$ 0.209
15% trimmed means reported (due to EIG instability);
SDR excluded due to limited scalability under uniform edge dropout.
In this section, we report the recovery performance and numerical efficiency of our propose DQGPM for SE(3) synchronization tasks on both synthetic and real data, where the real data experiments are implemented on multiple point-set registration problem. We also compare our method with two existing matrix-based methods, which are the spectral decomposition (EIG) described in Arrigoni et al. (2016b), and the semidefinite relaxation (SDR) approach introduced in Rosen et al. (2019). All experiments were performed in Python 3.9 on a computer running Windows 11, equipped with an AMD Ryzen 7 6800HS Creator Edition (3.20 GHz) processor and 16 GB of RAM. Experimental results on synthetic data with additive noise are presented in Section 4.2, while Section 4.3 reports results obtained from real data in the multiple point-set registration experiment.
@@ 4.1 Data Generation and Evaluation Metrics
Here, $\hat{\bm{x}}=(\hat{x}_{1},\ldots,\hat{x}_{n})\in\mathrm{UDQ}^{n}$ is the ground-truth pose vector, and term $\bm{\Xi}=(\xi_{ij})\in\mathrm{UDQ}^{n\times n}$ is a Hermitian dual quaternion noise matrix, all their entries are sampled i.i.d. as unit dual quaternions. The matrix $\bm{1}_{n\times n}$ denotes the all-one dual quaternion matrix. The term $\bm{\Xi}-\bm{1}_{n\times n}$ corresponds to the additive noise $\bm{\Delta}$ in the model (1.2).
@@
To represent elements in $\mathrm{SE}(3)$ for both the ground-truth poses and the noise terms, we sample $(\bm{u},\theta,\bm{t})$ to generate an unit dual quaternion $x$. Each sampled dual quaternion $x$ is constructed from a random rotation represented by $q\in\mathbb{H}$ and translation $\bm{t}\in\mathbb{R}^{3}$: we first form the unit quaternion $q$ from a randomly sampled axis $\bm{u}\in\mathbb{R}^{3}$ and angle $\theta\in[0,2\pi)$ via (B.1), and generate a randomly sampled $\bm{t}$ then convert $(q,\bm{t})$ to $x$ using (B.2).
@@
We follow the evaluation and gauge-alignment procedure of Hadi et al. (2024). Define ground-truth poses as $\hat{\bm{x}}=\{(\hat{q}_{i},\hat{t}_{i})\}_{i=1}^{n}\in\mathrm{SE}(3)^{n}$ and estimates as $\bm{x}=\{(q_{i},t_{i})\}_{i=1}^{n}\in\mathrm{SE}(3)^{n}$, where $\hat{q}_{i},q_{i}\in\mathbb{H}$ are unit quaternions and $\hat{t}_{i},t_{i}\in\mathbb{R}^{3}$ are translations. We compute the estimation error as the distance between $\bm{x}$} and the right-aligned ground-truth $\hat{\bm{x}}z$, where $z$ is the optimal aligner that minimizes the distance between $\bm{x}$ and $\hat{\bm{x}}z$. Specifically, we get the aligning SE(3) element $z=(q,t)$ by
@@
$d_{R}(q_{1},q_{2})$ $=2\arccos\bigl(2\langle q_{1},q_{2}\rangle^{2}-1\bigr),$
$d_{T}(t_{1},t_{2})$ $=\lVert t_{1}-t_{2}\rVert_{2},$
where $\langle\cdot,\cdot\rangle$ denotes the Euclidean inner product in $\mathbb{R}^{4}$. We report error_r and error_t as the mean of $d_{R}$ and $d_{T}$, respectively, over all $n$ targets.
4.2 Recovery Performance and Computational Time
For each ground-truth pose $\hat{x}_{i}$, we sample the rotation axis $\bm{u}$ uniformly from the unit sphere in $\mathbb{R}^{3}$ and the rotation angle independently and uniformly from $0,2\pi)$. The translation $\bm{t}=(t_{1},t_{2},t_{3})\in\mathbb{R}^{3}$ has i.i.d. entries following $\mathcal{N}(0,1)$. To generate the noise matrix, we again sample rotation axes uniformly from the unit sphere, while rotation angles (in degrees) are drawn i.i.d. from $\mathcal{N}(0,\sigma_{r}^{2})$ with $\sigma_{r}\in\{5^{\circ},10^{\circ},15^{\circ},20^{\circ}\}$. The corresponding translation noise has i.i.d. entries drawn from $\mathcal{N}(0,\sigma_{t}^{2})$ with $\sigma_{t}\in\{0.05,0.10,0.15,0.2\}$, respectively.
We run synthetic experiments with problem size $n=100$ under three observation rates
$p\in\{0.05,0.08,0.3\},$
corresponding to approximately $95\%$, $92\%$, and $70\%$ missing pairs, respectively. For each setting, we execute Algorithm 1 until the residual drops below $10^{-5}$, and then project the output elementwise onto $\mathrm{UDQ}^{n}$. All results are averaged over 100 independent trials.
Table 3: CPU time of DQGPM and EIG across noise levels and observation rates.
$(0.05,5^{\circ})$ $(0.10,10^{\circ})$ $(0.15,15^{\circ})$ $(0.20,20^{\circ})$
$p=0.05$ DQGPM 0.011 0.011 0.010 0.010
EIG 0.014 0.014 0.015 0.014
$p=0.08$ DQGPM 0.007 0.007 0.008 0.007
EIG 0.014 0.011 0.013 0.011
$p=0.30$ DQGPM 0.010 0.011 0.010 0.011
EIG 0.017 0.017 0.017 0.019
Table 2 reports the rotation and translation errors (error_r, error_t) of DQGPM and the baseline EIG across various noise levels and observation rates. The two metrics exhibit consistent trends, and DQGPM is uniformly more accurate and stable, with the largest gains in sparse graphs ($p=0.05$). DQGPM maintains small errors across all noise levels, whereas EIG incurs larger errors with substantially higher variance. As the graph becomes denser ($p=0.08$ and $p=0.30$), both methods improve and the performance gap narrows; nonetheless, DQGPM continues to achieve lower errors and reduced variance, retaining a mild but consistent advantage in the dense regime. Besides, we record the total CPU time consumed by each method in Table 3. It can be observed that DQGPM is consistently faster than EIG across all settings, and its runtime is insensitive to the noise level.
((a)) Rotation error
((b)) Translation error
4.3 Efficiency and Accuracy on Real Data
We evaluate our method on real data from the Stanford 3D Scanning Repository [^1] http://graphics.stanford.edu/data/3Dscanrep and compare it against EIG and SDR methods. To construct the measurement matrix, we first use Iterative Closest Point (ICP) algorithm (Besl and McKay, 1992) to estimate relative rigid motions, on which we apply additional rigid-motion perturbation. Specifically, the rotation angle (in degrees) is sampled from $\mathcal{N}(0,1)$, the rotation axis is drawn uniformly from the unit sphere, and the translation vector (in meters) has i.i.d. entries following $\mathcal{N}(0,0.001^{2})$, consistent with the scale of translations in the datasets. All results are averaged over 100 independent trials.
((a)) Bunny
((b)) Buddha
((c)) Dragon
((d)) Armadillo
Table 4: Performance on four real datasets under $(\sigma_{t},\sigma_{r})=(0.001,1^{\circ})$: missing rate, time, and errors.
@@ 4.3 Efficiency and Accuracy on Real Data
Missing 48.00% 13.33% 66.67% 18.67% 60.44% 19.44% 58.33%
Time DQGPM 0.0072 0.0023 0.0216 0.0024 0.0126 0.0027 0.0137
SDR 0.0954 0.0887 0.1078 0.0937 0.1226 0.1020 0.1114
EIG 0.0030 0.0032 0.0036 0.0032 0.0051 0.0053 0.0071
Error_r DQGPM 0.0188 0.0192 0.0233 0.0218 0.0249 0.0198 0.0198
SDR 3.6349 0.1395 3.1227 0.2079 3.1307 0.5665 3.1251
EIG 3.4519 0.0923 3.1022 0.1704 3.1346 0.3637 3.1273
Error_t DQGPM 0.0012 0.0007 0.0016 0.0009 0.0019 0.0012 0.0013
SDR 0.0675 0.0014 0.0003 0.0014 0.0003 0.0013 0.0003
EIG 0.0506 0.0006 0.0002 0.0009 0.0003 0.0006 0.0003
Table 5: Cross-sectional views of the aligned point clouds on four real datasets for different methods.
Dataset DQGPM SDR EIG
Buddha
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Dragon
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Armadillo
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]

### v2

@@ 3 Iterative Estimation Performance
It is worth noting the improvement in estimation accuracy provided by DQGPM over the rounded spectral estimator. Theorem 3.2 establishes that, asymptotically as $k\to\infty$ (assuming convergent iterates and a small noise constant regime where $\|\bm{\Delta}\|_{\mathrm{op},\mathrm{st}}\leq an$ for sufficiently small $a>0$), the DQGPM estimation error satisfies:
$\operatorname{d}_{\mathrm{st}}(\bm{x}^{\infty},\hat{\bm{x}})\lesssim\frac{12}{n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,\mathrm{st}},$
$|\operatorname{d}_{\mathcal{I}}(\bm{x}^{\infty},\hat{\bm{x}})|\leq\delta^{\infty}_{\mathcal{I}}\lesssim\frac{396B+124\gamma}{45n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,\mathrm{st}}.$
In contrast, Theorem 2.8 guarantees an error bound for the rounded spectral estimator of order $8\|\bm{\Delta}\|_{\mathrm{op}}/\sqrt{n}$. Comparing these rates, DQGPM yields a tighter bound provided that $\|\bm{\Delta}\hat{\bm{x}}\|_{2,\mathrm{st}}\leq c\sqrt{n}\|\bm{\Delta}\|_{\mathrm{op},\mathrm{st}}$ for a constant $c<2/3$. Note that the trivial bound $c=1$ holds automatically by definition. In many stochastic noise models, however, $\hat{\bm{x}}$ is not aligned with the top singular directions of $\bm{\Delta}$, so $\|\bm{\Delta}\hat{\bm{x}}\|_{2}$ is typically much smaller than the worst-case bound (e.g., $c\approx 1/2$ for standard i.i.d. Gaussian noise), which highlights the advantage of the DQGPM refinement.
Table 2: Rotation and translation errors in synthetic experiments of DQGPM, SPEC, and EIG across noise levels and observation rates.
$p=0.05$ Error_r DQGPM 0.132 $\pm$ 0.042 0.230 $\pm$ 0.048 0.326 $\pm$ 0.051 0.424 $\pm$ 0.060
SPEC 1.639 $\pm$ 1.971 1.889 $\pm$ 2.022 1.854 $\pm$ 1.862 2.035 $\pm$ 1.823
EIG 0.174 $\pm$ 0.156 0.270 $\pm$ 0.099 0.406 $\pm$ 0.190 0.585 $\pm$ 0.572
Error_t DQGPM 0.102 $\pm$ 0.032 0.190 $\pm$ 0.046 0.279 $\pm$ 0.061 0.369 $\pm$ 0.078
SPEC 0.480 $\pm$ 0.530 0.532 $\pm$ 0.504 0.590 $\pm$ 0.472 0.660 $\pm$ 0.455
EIG 0.551 $\pm$ 1.109 1.609 $\pm$ 6.676 0.899 $\pm$ 1.478 1.043 $\pm$ 1.359
$p=0.30$ Error_r DQGPM 0.027 $\pm$ 0.001 0.054 $\pm$ 0.002 0.082 $\pm$ 0.004 0.111 $\pm$ 0.005
SPEC 0.032 $\pm$ 0.013 0.192 $\pm$ 0.867 0.094 $\pm$ 0.037 0.141 $\pm$ 0.126
EIG 0.098 $\pm$ 0.618 0.134 $\pm$ 0.614 0.106 $\pm$ 0.021 0.219 $\pm$ 0.613
Error_t DQGPM 0.021 $\pm$ 0.001 0.042 $\pm$ 0.002 0.063 $\pm$ 0.003 0.085 $\pm$ 0.005
SPEC 0.023 $\pm$ 0.001 0.045 $\pm$ 0.002 0.067 $\pm$ 0.004 0.090 $\pm$ 0.005
EIG 0.137 $\pm$ 0.441 0.415 $\pm$ 2.123 0.249 $\pm$ 0.677 0.194 $\pm$ 0.257
SDR excluded due to limited scalability under uniform edge dropout.
@@
Suppose the assumptions in Theorem 3.2 hold. Let $\bm{C}\in\mathbb{DH}^{n\times n}$ be Hermitian and admits a unique dominant eigenvalue with multiplicity one. Denote its right eigenvalues by $\lambda_{1}>\lambda_{2}\geq\cdots\geq\lambda_{n}$, where each $\lambda_{i}=\lambda_{i,\mathrm{st}}+\lambda_{i,\mathcal{I}}\epsilon$ is a dual number. If we run Algorithm 1 for $K_{\mathrm{init}}$ iterations satisfying
$K_{\mathrm{init}}\geq$ $\log_{r}\left({\frac{70|\alpha_{2,\mathrm{st}}|+69M_{\mathrm{st}}}{|\alpha_{1,\mathrm{st}}|}}\right),$
then, initializing Algorithm 2 with $\bm{x}^{0}:=\Pi(\sqrt{n}\bm{w}^{K_{\mathrm{init}}})$, the resulting sequence $\{\bm{x}^{k}\}_{k\geq 0}$ satisfies (6) and (10). Here $r:=\bigl|\lambda_{1,_{\mathrm{st}}}/\lambda_{2,_{\mathrm{st}}}\bigr|>1$, $M_{{}_{\mathrm{st}}}$ is a real constant depending on $\{\alpha_{j},\lambda_{j}\}_{j=1}^{n}$, where $\alpha_{i}$ is the projection of initializer $\bm{w}^{0}$ onto the $i$-th orthonormal eigenvector.
@@ 4 Numerical Experiments
In this section, we report the recovery performance and numerical efficiency of our proposed DQGPM for SE(3) synchronization tasks on both synthetic and real data, where the real data experiments are implemented on multiple point-set registration problem. We also compare our method with three existing matrix-based methods, which are the spectral decomposition (EIG) described in Arrigoni et al. (2016b), the semidefinite relaxation approach (SDR) introduced in Rosen et al. (2019) and the recent spectral relaxation method (SPEC) proposed in Doherty et al. (2022). All experiments were performed in Python 3.9 on a computer running Windows 11, equipped with an AMD Ryzen 7 6800HS Creator Edition (3.20 GHz) processor and 16 GB of RAM. Experimental results on synthetic data with additive noise are presented in Section 4.2, while Section 4.3 reports results obtained from real data in the multiple point-set registration experiment. The implementation code is available at https://github.com/jnzhao333/dq_sync.
Table 3: Cross-sectional views of the aligned point clouds on three real datasets for different methods.
Dataset DQGPM SPEC SDR EIG
Buddha
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Dragon
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Armadillo
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ 4.1 Data Generation and Evaluation Metrics
Here, $\hat{\bm{x}}=(\hat{x}_{1},\ldots,\hat{x}_{n})\in\mathrm{UDQ}^{n}$ is the ground truth pose vector, and term $\bm{\Xi}=(\xi_{ij})\in\mathrm{UDQ}^{n\times n}$ is a Hermitian dual quaternion noise matrix, all their entries are sampled i.i.d. as unit dual quaternions. The matrix $\bm{1}_{n\times n}$ denotes the all-one dual quaternion matrix. The term $\bm{\Xi}-\bm{1}_{n\times n}$ corresponds to the additive noise $\bm{\Delta}$ in the model (2).
@@
To represent elements in $\mathrm{SE}(3)$ for both the ground truth poses and the noise terms, we sample $(\bm{u},\theta,\bm{t})$ to generate a unit dual quaternion $x$. Each sampled dual quaternion $x$ is constructed from a random rotation represented by $q\in\mathbb{H}$ and translation $\bm{t}\in\mathbb{R}^{3}$: we first form the unit quaternion $q$ from a randomly sampled axis $\bm{u}\in\mathbb{R}^{3}$ and angle $\theta\in[0,2\pi)$ via (11), and generate a randomly sampled $\bm{t}$ then convert $(q,\bm{t})$ to $x$ using (12).
@@
We follow the evaluation and gauge-alignment procedure of Hadi et al. (2024). Define ground truth poses as $\hat{\bm{x}}=\{(\hat{q}_{i},\hat{t}_{i})\}_{i=1}^{n}\in\mathrm{SE}(3)^{n}$ and estimates as $\bm{x}=\{(q_{i},t_{i})\}_{i=1}^{n}\in\mathrm{SE}(3)^{n}$, where $\hat{q}_{i},q_{i}\in\mathbb{H}$ are unit quaternions and $\hat{t}_{i},t_{i}\in\mathbb{R}^{3}$ are translations. We compute the estimation error as the distance between $\bm{x}$} and the right-aligned ground truth $\hat{\bm{x}}z$, where $z$ is the optimal aligner that minimizes the distance between $\bm{x}$ and $\hat{\bm{x}}z$. Specifically, we get the aligning SE(3) element $z=(q,t)$ by
@@
$\operatorname{d}_{R}(q_{1},q_{2})$ $=2\arccos\bigl(2\langle q_{1},q_{2}\rangle^{2}-1\bigr),$
$\operatorname{d}_{T}(t_{1},t_{2})$ $=\lVert t_{1}-t_{2}\rVert_{2},$
where $\langle\cdot,\cdot\rangle$ denotes the Euclidean inner product in $\mathbb{R}^{4}$. We report error_r and error_t as the mean of $\operatorname{d}_{R}$ and $\operatorname{d}_{T}$, respectively, over all $n$ targets.
Table 4: Performance on four real datasets under $(\sigma_{t},\sigma_{r})=(0.01,10^{\circ})$: missing rate, time, and errors.
@@
Missing 48.00% 18.67% 66.67% 19.44% 60.44% 19.44% 58.33%
Time DQGPM 0.0010 0.0003 0.0021 0.0003 0.0014 0.0003 0.0009
SPEC 0.0018 0.0043 0.0041 0.0043 0.0042 0.0028 0.0028
SDR 0.0954 0.2166 0.1893 0.2306 0.1964 0.2410 0.1907
EIG 0.0006 0.0008 0.0007 0.0009 0.0007 0.0007 0.0006
Error_r DQGPM 0.1631 0.0898 0.2118 0.0935 0.1648 0.1095 0.1761
SPEC 3.4489 0.1826 3.1018 0.2404 3.1359 0.5723 3.1249
SDR 3.6349 0.1937 3.1181 0.2843 3.1320 0.5757 3.1236
EIG 3.4460 0.1736 3.0993 0.2623 3.1349 0.5785 3.1260
Error_t DQGPM 0.0107 0.0044 0.0117 0.0047 0.0085 0.0056 0.0088
SPEC 0.0656 0.0038 0.0021 0.0038 0.0024 0.0040 0.0026
SDR 0.0675 0.0042 0.0021 0.0042 0.0024 0.0043 0.0027
EIG 0.0509 0.0039 0.0021 0.0039 0.0024 0.0042 0.0026
4.2 Recovery Performance and Computational Time
For each ground truth pose $\hat{x}_{i}$, we sample the rotation axis $\bm{u}$ uniformly from the unit sphere in $\mathbb{R}^{3}$ and the rotation angle independently and uniformly from $0,2\pi)$. The translation $\bm{t}=(t_{1},t_{2},t_{3})\in\mathbb{R}^{3}$ has i.i.d. entries following $\mathcal{N}(0,1)$. To generate the noise matrix, we again sample rotation axes uniformly from the unit sphere, while rotation angles (in degrees) are drawn i.i.d. from $\mathcal{N}(0,\sigma_{r}^{2})$ with $\sigma_{r}\in\{5^{\circ},10^{\circ},15^{\circ},20^{\circ}\}$. The corresponding translation noise has i.i.d. entries drawn from $\mathcal{N}(0,\sigma_{t}^{2})$ with $\sigma_{t}\in\{0.05,0.10,0.15,0.2\}$, respectively.
We run synthetic experiments with problem size $n=100$ under three observation rates $p\in\{0.05,0.3\}$, corresponding to approximately $95\%$ and $70\%$ missing pairs, respectively. For each setting, we execute Algorithm 1 until the residual drops below $10^{-5}$, and then project the output elementwise onto $\mathrm{UDQ}^{n}$. All results are averaged over 100 independent trials.
Table 5: CPU time of DQGPM, SPEC and EIG across noise levels and observation rates.
$(0.05,5^{\circ})$ $(0.10,10^{\circ})$ $(0.15,15^{\circ})$ $(0.20,20^{\circ})$
$p=0.05$ DQGPM 0.005 0.005 0.005 0.004
SPEC 0.084 0.089 0.069 0.069
EIG 0.006 0.005 0.005 0.004
$p=0.30$ DQGPM 0.002 0.001 0.001 0.001
SPEC 0.143 0.140 0.140 0.142
EIG 0.003 0.004 0.004 0.004
Table 2 reports the rotation and translation errors (error_r, error_t) of DQGPM and the baselines SPEC and EIG across various noise levels and observation rates. The two metrics exhibit consistent trends, and DQGPM is uniformly more accurate and stable, with the largest gains in sparse graphs ($p=0.05$). DQGPM maintains small errors across all noise levels, whereas EIG incurs larger errors with substantially higher variance. As the graph becomes denser ($p=0.30$), both methods improve and the performance gap narrows; nonetheless, DQGPM continues to achieve lower errors and reduced variance, retaining a mild but consistent advantage in the dense regime. Besides, we record the total CPU time consumed by each method in Table 5. It can be observed that DQGPM is consistently faster than EIG across all settings, and its runtime is insensitive to the noise level.
(a) Rotation error
(b) Translation error
4.3 Real Data: Stanford Bunny and 3DMatch
We evaluate our method on real data from the Stanford 3D Scanning Repository [^1] http://graphics.stanford.edu/data/3Dscanrep and 3DMatch [^2] https://3dmatch.cs.princeton.edu/#geometric-registration-benchmark dataset, and compare it against EIG, SDR and SPEC methods. To construct the measurement matrix, we first use Iterative Closest Point (ICP) algorithm (Besl and McKay, 1992) to estimate relative rigid motions, on which we apply additional rigid-motion perturbation. Specifically, the rotation angle (in degrees) is sampled from $\mathcal{N}(0,10^{2})$, the rotation axis is drawn uniformly from the unit sphere, and the translation vector (in meters) has i.i.d. entries following $\mathcal{N}(0,0.01^{2})$, consistent with the scale of translations in the datasets. All results are averaged over 100 independent trials.
(a) Bunny
(b) Buddha
(c) Dragon
(d) Armadillo